# Dynamical phase transitions in the non-reciprocal Ising model

### v1

@@ II.4 Classical continuous time crystals
where $\eta(t)$ is a white noise with $\langle\eta\rangle=0$ and $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t-t^{\prime})$. Here, $\phi$ can be seen as the argument of a complex number $z(t)=Ae^{i\phi(t)}$ whose amplitude $A$ is very quickly relaxing to its steady-state value. In this very simple system, we observe that the average value $\braket{\phi}$ always oscillates (averaging the equation of motion leads to $\braket{\phi(t)}=\braket{\phi(0)}+\omega t$). However, this oscillator only keeps track of time (i.e. of its initial phase) for so long. To see that, consider the correlation function $C(\tau)=\langle e^{i[\phi(t)-\phi(0)]}\rangle$ which describes the average phase that an assembly of independent clocks would have at time $t$, provided that they were synchronized at time $t=0$. Here, it behaves as 333 This can be seen by integrating $d\phi_{t}=\omega dt+\sqrt{2D_{\phi}}dW_{t}$ where $W_{t}$ is a standard Wiener process, leading to $\phi(t)-\phi(0)=\omega t+\sqrt{2D_{\phi}}W_{t}$. We find $C(\tau)=e^{i\omega t}\langle e^{i\sqrt{2D_{\phi}}W_{t}}\rangle$. As $W_{t}$ follows a normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=t)$, its moment distribution function is $M_{W_{t}}(x)\equiv\mathbb{E}(e^{xW_{t}})=e^{\mu x+\sigma^{2}x^{2}/2}$ from which the result is found.
$C(\tau)\propto\cos(\omega t)e^{-t/\tau}$ (6)
where $\tau=1/D_{\phi}$, meaning that the oscillator looses track of its original phase after $\mathcal{Q}=\omega\tau$ oscillations ($\mathcal{Q}$ is the quality factor of the oscillator). In fact, any noisy oscillator behaves in this way at long times.
Thermodynamic bounds on the number $\mathcal{Q}$ of coherent oscillations have been considered [124, 125, 126, 127], and suggest that an infinite amount of energy (more precisely of entropy production) is needed to obtain infinite coherence. It is however not excluded that a macroscopic system in the thermodynamic limit can have an infinite coherence time with a finite cost per unit volume. Our numerical results suggest that this indeed happens in our model: we evaluate the coherence time of oscillations, which scales as $\tau(L)\sim L^{3}$ (in $d=3$). This points to the existence of perfectly coherent oscillations in the thermodynamic limit, in which the rate of entropy production per unit volume is finite. In this sense, our model can be seen as a classical continuous time crystal [128, 129, 130, 131, 132, 133], see also Refs. [128, 134, 135, 136, 129, 137, 138, 139] on time crystals. We conjecture that this property may extend to other systems with locally coupled oscillators that spontaneously break a continuous time-translation symmetry.
@@ III Outline and summary
The outline of the paper is as follows: we present the non-reciprocal Ising model with general asymmetric coupling between the two species and derive its mean-field equation in Sec. IV. In Secs. V and  VIIXI we restrict ourselves to the fully anti-symmetric case in which the system exhibits $C_{4}$ symmetry. In Sec. IV we show that the mean-field equation predicts three distinct phases: disorder, time-dependent order that we dub “swap”, and static order, and analyze the bifurcations that separate the phases. Details on the Monte-Carlo simulations are give in Sec. VI. Our numerical simulations in 2D are presented in Sec. VII, suggesting that the swap phase is unstable in 2D due to spiral defects. In Sec. VIII we show that the swap phase is stable in 3D, and calculate the static critical exponents of the phase transition from disorder to swap. These critical exponents are in excellent agreement with the equilibrium XY critical exponents, which is compatible with renormalization-group studies and with the emergent continuous time-translation symmetry. The coherence time of the magnetization is shown to grow with system size suggesting a time-crystal behavior in the thermodynamic limit. In Sec. IX, we provide numerical evidence and theoretical arguments that suggest that the static-order phase is destroyed in any finite dimension due to droplet growth. In Sec. X we discuss how the system coarsens into the swap and static-order phases. In Sec. XII we allow the presence of inter-species reciprocal couplings, which reduces the symmetry to $\mathbb{Z}_{2}$, and accounts for a general asymmetric interaction. We present mean-field analysis and 2D and 3D simulation results. These results reveal that the swap phase stability is unchanged with respect to the fully asymmetric case (i.e. it is stable in 3D but not in 2D). However, static order can be retrieved by an intriguing droplet-capture mechanism. Contrary to the mean-field prediction, we find no evidence of a direct transition between the swap and the static-order phase. We discuss our results and draw conclusions in Sec. XIII.
@@ IV The model
In the equilibrium Ising model, adjacent on-lattice spins ($\sigma=\pm 1$) stochastically align with one another through an energy function. Its dynamics is given by kinetic Ising models which, while lacking uniqueness, always operate by minimizing the global free energy. In what follows, we generalize these kinetic Ising models to the case of non-reciprocal interactions where such global potential function does not exist. To this aim, we define a Glauber dynamics [140] where each spin minimizes its own selfish energy. Similar formulations were used among the physical sciences [65, 52, 35, 141, 142, 143, 20, 144, 145, 146, 6].
IV.1 The non-reciprocal (NR) Ising model
We consider two spin species, labeled by Greek indices $\alpha=A,B$, located on a $d$-dimensional cubic lattice of linear size $L$. Each site $i\in\{1,...,L^{d}\}$ contains one A-spin and one B-spin (Fig. 1, left).
The state of the system is described by a vector
@@ IV.1 The non-reciprocal (NR) Ising model
We consider transition rates given by the generalized Glauber dynamics [140]
@@
where $\tau$ is the single spin-flip time-scale, $T$ is the temperature, $k_{B}$ the Boltzmann constant and
@@
is the difference in selfish energy between the states. Here $F^{\alpha}_{i}$ is a single-spin-flipping operator, i.e., when acted on $\vec{\sigma}$ it flips the $\alpha$-species spin on site $i$ from $\sigma^{\alpha}_{i}$ to $-\sigma^{\alpha}_{i}$ while keeping all the other spins unchanged. For instance,
@@
Finally, $w(\vec{\sigma}|\vec{\sigma}^{\prime})$ is the transition rate from configuration $\vec{\sigma}$ to $\vec{\sigma}^{\prime}$.
The selfish energy is
@@
and $J,K_{+},K_{-}>0$. The sum in Eq. (11) runs over nearest neighbors of $i$ and summation over spin indices $\beta$ is implied. The selfish energy indicates that spins of the same species tend to align with their neighbors, while spins of different species interact non-reciprocally (see Fig. 1, right): spins $A$ tend to align with spins $B$, whereas spins $B$ either tend to align with spins $B$ but to a lesser extent (if $K_{+}>K_{-}$), or tend to anti-align with spins $A$ (if $K_{+}<K_{-}$). The latter case will be the main focus of this paper as it leads to a rich phenomenology related to limit cycles. For simplicity, note that inter-species interactions only occur on the same site.
Finally, in the continuous-time limit, the probability $P(\vec{\sigma},t)$ to observe a specific configuration of spins $\vec{\sigma}$ at time $t$ evolves according to the master equation
@@
Fig. 1: Schematic drawing of the non-reciprocal (NR) Ising model. The model includes two species per site with two interaction types: (i) reciprocal nearest-neighbors interaction of strength $J$ between same-species spins and (ii) non-reciprocal same-site interaction of strength $K_{+}+K_{-}$ for species A (blue) and $K_{+}-K_{-}$ for species B (red), between spins of different species.
Fig. 2: Schematic drawing of the mapping between the state of two spins located on the same site ($\sigma^{A}$ and $\sigma^{B}$) into its corresponding angle variable $\theta$. The color associated to each quadrant is used throughout the paper to indicate the corresponding spin state. The direction of the cycle is determined by the sign of $K_{-}$ (throughout the paper $K_{-}>0$ which implies clockwise direction).
@@ IV.2 Broken detailed balance
A Markov chain with rates $w$ satisfies detailed balance if it has a stationary distribution $P_{\text{s}}$ such that
$\frac{w(\vec{\sigma}\to\vec{\sigma}^{\prime})}{w(\vec{\sigma}^{\prime}\to\vec{\sigma})}=\frac{P_{\text{s}}(\vec{\sigma}^{\prime})}{P_{\text{s}}(\vec{\sigma})}.$ (14)
Let us assume that our kinetic model possesses an (unknown) stationary distribution $P_{\text{s}}(\vec{\sigma})$. Considering a loop $\vec{\sigma}_{1}\to\vec{\sigma}_{2}\to\cdots\to\vec{\sigma}_{n}$ in state space, as well as the same loop travelled in opposite direction, detailed balance implies [147]
@@
$\!\!\!\frac{w\left(\uparrow\uparrow\to\uparrow\downarrow\right)w\left(\uparrow\downarrow\to\downarrow\downarrow\right)w\left(\downarrow\downarrow\to\downarrow\uparrow\right)w\left(\downarrow\uparrow\to\uparrow\uparrow\right)}{w\left(\uparrow\downarrow\to\uparrow\uparrow\right)w\left(\downarrow\downarrow\to\uparrow\downarrow\right)w\left(\downarrow\uparrow\to\downarrow\downarrow\right)w\left(\uparrow\uparrow\to\downarrow\uparrow\right)}=e^{8\tilde{K}_{-}}$ (17)
which means that detailed balance is broken in our model whenever $\tilde{K}_{-}\neq 0$.
@@ IV.3 The mean-field (MF) equation
The average magnetization of a single spin at time $t$ is given by
$\langle\sigma^{\alpha}_{i}\rangle\left(t\right)=\sum_{\vec{\sigma}}\sigma^{\alpha}_{i}P({\vec{\sigma}},t).$ (18)
Using Eq. (13) and changing summation variables we get
@@
Using Eq. (8) and substituting the selfish energy (Eq. (11)) we obtain
@@
where we have further used the identity $x\tan(xb)=\tan(b)$ for $x=\pm 1$. Applying the mean-field (MF) approximation, $\langle g(\vec{\sigma})\rangle=g(\langle\vec{\sigma}\rangle)$ and defining the average magnetization as $\langle\sigma^{\alpha}_{i}\rangle=m^{\alpha}_{i}$ yields
$\partial_{t}m_{i}^{\alpha}=-\frac{m_{i}^{\alpha}}{\tau}+\frac{1}{\tau}\tanh\left(\frac{J\sum_{j\,{\rm nn\,of}\,i}m_{j}^{\alpha}+K_{\alpha\beta}m_{i}^{\beta}}{k_{B}T}\right).$ (21)
@@
where $D\equiv J/(k_{B}T)$ and $\tilde{K}_{\alpha\beta}\equiv K_{\alpha\beta}/(k_{B}T)$. Equation (23) is the MF equation of the general NR Ising model. Note that it is invariant under the $\mathbb{Z}_{2}$ symmetry $(m_{A},m_{B})\mapsto(-m_{A},-m_{B})$.
@@ V The fully anti-symmetric case
In this section as well as in Secs. VIIXI, we will study the fully anti-symmetric case, in which $K_{+}=0$. To simplify the notations, we write $K_{-}\equiv K$ (and $\tilde{K}_{-}\equiv\tilde{K}$). In this case, the MF equation reads
@@
where $\varepsilon_{\alpha\beta}$ is the Levi-Civita symbol. Equation (24) has the additional symmetry $(m_{A},m_{B})\mapsto(-m_{B},m_{A})$, so overall it is invariant under $C_{4}$.

### v2

@@ II.4 Classical continuous time crystals
where $\eta(t)$ is a white noise with $\langle\eta\rangle=0$ and $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t-t^{\prime})$. Here, $\phi$ can be seen as the argument of a complex number $z(t)=Ae^{i\phi(t)}$ whose amplitude $A$ is very quickly relaxing to its steady-state value. In this simple system, we observe that the average value $\braket{\phi}$ always oscillates (averaging the equation of motion leads to $\braket{\phi(t)}=\phi(0)+\omega t$). However, this oscillator only keeps track of time (i.e. of its initial phase) for so long. To see that, consider the correlation function $C(t)=\langle e^{i[\phi(t)-\phi(0)]}\rangle$ which describes the average phase that an assembly of independent clocks would have at time $t$, provided that they were synchronized at time $t=0$. Here, it behaves as 333 This can be seen by integrating $d\phi_{t}=\omega dt+\sqrt{2D_{\phi}}dW_{t}$ where $W_{t}$ is a standard Wiener process, leading to $\phi(t)-\phi(0)=\omega t+\sqrt{2D_{\phi}}W_{t}$. We find $C(t)=e^{i\omega t}\langle e^{i\sqrt{2D_{\phi}}W_{t}}\rangle$. As $W_{t}$ follows a normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=t)$, its moment generating function is $M_{W_{t}}(x)\equiv\mathbb{E}(e^{xW_{t}})=e^{\mu x+\sigma^{2}x^{2}/2}$ from which the result is found.
$C(t)\propto\cos(\omega t)e^{-t/\tau}$ (6)
where $\tau=1/D_{\phi}$, meaning that the oscillator loses track of its original phase after $\mathcal{Q}=\omega\tau$ oscillations ($\mathcal{Q}$ is known as the quality factor of the noisy oscillator). In fact, any noisy oscillator behaves in this way at long times [129].
Thermodynamic bounds on the number $\mathcal{Q}$ of coherent oscillations have been considered [130, 131, 132, 133], and suggest that an infinite amount of energy (more precisely of entropy production) is needed to obtain infinite coherence. It is however not excluded that a macroscopic system in the thermodynamic limit can have an infinite coherence time with a finite cost per unit volume. Our numerical results suggest that this indeed happens in our model: we evaluate the coherence time of oscillations, which scales as $\tau(L)\sim L^{3}$ (in $d=3$). This points to the existence of perfectly coherent oscillations in the thermodynamic limit. Additionally, we measure the entropy production rate per unit volume and find that it converges to a finite value, independent of system size. In this sense, our model can be seen as a classical continuous time crystal [134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147]. We conjecture that this property may extend to other systems with locally coupled oscillators that spontaneously break a continuous time-translation symmetry.
@@ III Outline and summary
The outline of the paper is as follows: we present the nonreciprocal Ising model with general asymmetric coupling between the two species and derive its mean-field equation in Sec. IV. In Secs. V and  VIIXI we restrict ourselves to the fully anti-symmetric case in which the system exhibits $C_{4}$ symmetry. In Sec. IV we show that the mean-field equation predicts three distinct phases: disorder, time-dependent order that we dub “swap”, and static order, and analyze the bifurcations that separate the phases. Details on the Monte-Carlo simulations are given in Sec. VI. Our numerical simulations in 2D are presented in Sec. VII, suggesting that the swap phase is unstable in 2D due to spiral defects. In Sec. VIII we show that the swap phase is stable in 3D, and calculate the static critical exponents of the phase transition from disorder to swap. These critical exponents are in excellent agreement with the equilibrium XY critical exponents, which is compatible with renormalization-group studies and with the breaking of a continuous time-translation symmetry. We calculate the coherence time of the magnetization and show that it grows with system size, suggesting a time-crystal behavior in the thermodynamic limit. In Sec. IX, we provide numerical evidence and theoretical arguments that suggest that the static-order phase is destroyed in any finite dimension due to droplet growth. In Sec. X we discuss how the system coarsens into swap and static-order states. The phase diagram of the fully anti-symmetric case is summarized in Sec. XI. In Sec. XII we allow the presence of inter-species reciprocal couplings, which reduces the symmetry to $\mathbb{Z}_{2}$, and accounts for a general asymmetric interaction. We present mean-field analysis and 2D and 3D simulation results. These results reveal that the swap phase stability is unchanged with respect to the fully anti-symmetric case (i.e. it is stable in 3D but not in 2D). However, static order can be restored by an intriguing droplet-capture mechanism. Contrary to the mean-field prediction, we find no evidence of a direct transition between the swap and the static-order phase. We summarize our results and conclusions in Sec. XIII.
@@ IV The model
In the equilibrium Ising model, adjacent on-lattice spins (${\sigma=\pm 1}$) stochastically align with one another through an energy function. Its dynamics is given by kinetic Ising models which, while lacking uniqueness, always operate by minimizing the global free energy. In what follows, we generalize these kinetic Ising models to the case of nonreciprocal interactions where such global potential function does not exist. To this aim, we define a Glauber dynamics [148] where each spin tends to minimize its own selfish energy. Similar formulations were used among the physical sciences [65, 52, 35, 149, 150, 151, 20, 152, 153, 154, 6].
IV.1 The nonreciprocal (NR) Ising model
We consider two spin species, labeled by Greek indices $\alpha=A,B$, located on a $d$-dimensional cubic lattice of linear size $L$. Each site $i\in\{1,...,L^{d}\}$ contains one A-spin and one B-spin (Fig. 1, left). The state of the system is described by a vector
@@ IV.1 The nonreciprocal (NR) Ising model
Fig. 1: Nonreciprocal (NR) Ising model. The model includes two species per site with two interaction types: (i) reciprocal nearest-neighbors interaction of strength $J$ between same-species spins and (ii) nonreciprocal same-site interaction of strength $K_{+}+K_{-}$ for species A (blue) and $K_{+}-K_{-}$ for species B (red), between spins of different species.
We consider transition rates given by the generalized Glauber dynamics [148] in which the transition rate $w(\vec{\sigma}|\vec{\sigma}^{\prime})$ from configuration $\vec{\sigma}$ to $\vec{\sigma}^{\prime}$ is given by
@@
where $T$ is the temperature, $k_{B}$ the Boltzmann constant, and $\tau$ is a characteristic time for single spin flips. In addition,
@@
is the difference in selfish energy between the states. Here $F^{\alpha}_{i}$ is a single-spin-flipping operator, i.e., when acting on $\vec{\sigma}$ it flips the $\alpha$-species spin on site $i$ from $\sigma^{\alpha}_{i}$ to $-\sigma^{\alpha}_{i}$ while keeping all the other spins unchanged. For instance,
@@
For the selfish energy, we choose
@@
and $J,K_{+},K_{-}>0$. The sum in Eq. (11) runs over nearest neighbors of $i$ and summation over spin indices $\beta$ is implied. The selfish energy indicates that spins of the same species tend to align with their neighbors, while spins of different species interact nonreciprocally (see Fig. 1, right): spins $A$ tend to align with spins $B$, whereas spins $B$ either tend to align with spins $B$ but to a lesser extent (if $K_{+}>K_{-}$), or tend to anti-align with spins $A$ (if $K_{+}<K_{-}$). The latter case will be the main focus of this paper as it leads to a rich phenomenology related to limit cycles. Note that in our model, inter-species interactions only occur between spins located on the same site.
In the continuous-time limit, the probability $P(\vec{\sigma},t)$ to observe a specific configuration of spins $\vec{\sigma}$ at time $t$ evolves according to the master equation
@@
As detailed in Sec. VI, we perform Monte-Carlo simulations that sample trajectories obtained from a discrete-time version of Eq. (13).
@@
Fig. 2: Mapping between the state of two spins located on the same site ($\sigma^{A}$ and $\sigma^{B}$) and their corresponding angle variable $\theta$. The color associated to each quadrant is used throughout the paper to indicate the corresponding spin state. The direction of the cycle is determined by the sign of $K_{-}$ (throughout the paper $K_{-}>0$ which implies clockwise direction).
@@ IV.2 Broken detailed balance
We now show that the nonreciprocal Ising model described in the previous section is out of equilibrium, in the sense that it breaks detailed balance.
A Markov chain with rates $w$ satisfies detailed balance if it has a stationary distribution $P^{\text{ss}}$ such that
$\frac{w(\vec{\sigma}\to\vec{\sigma}^{\prime})}{w(\vec{\sigma}^{\prime}\to\vec{\sigma})}=\frac{P^{\text{ss}}(\vec{\sigma}^{\prime})}{P^{\text{ss}}(\vec{\sigma})}.$ (14)
Let us assume that our kinetic model possesses an (unknown) stationary distribution $P^{\text{ss}}(\vec{\sigma})$. Considering a loop $\vec{\sigma}_{1}\to\vec{\sigma}_{2}\to\cdots\to\vec{\sigma}_{n}$ in state space, as well as the same loop travelled in opposite direction, detailed balance implies [155]
@@
$\!\!\!\frac{w\left(\uparrow\uparrow\to\uparrow\downarrow\right)w\left(\uparrow\downarrow\to\downarrow\downarrow\right)w\left(\downarrow\downarrow\to\downarrow\uparrow\right)w\left(\downarrow\uparrow\to\uparrow\uparrow\right)}{w\left(\uparrow\uparrow\to\downarrow\uparrow\right)w\left(\downarrow\uparrow\to\downarrow\downarrow\right)w\left(\downarrow\downarrow\to\uparrow\downarrow\right)w\left(\uparrow\downarrow\to\uparrow\uparrow\right)}=e^{8\tilde{K}_{-}}$ (17)
which means that detailed balance is broken in our model whenever $\tilde{K}_{-}\neq 0$. The breaking of detailed balance implies the system is driven away from equilibrium. Under certain assumptions on the physical realization of the model that go by the name of local detailed balance [156, 157, 158], the cycle produces entropy $\Delta S/k_{B}=8\tilde{K}_{-}$ (see Appendix A for details). We refer to Refs. [159, 160, 161, 162, 163] for discussions about irreversibility and entropy production in nonequilibrium phase transitions.
@@ IV.3 The mean-field (MF) equation
We now derive mean-field (MF) equations of motion for the magnetizations of each species from the master equation (13) (see Refs. [164, 165, 166]). The mean-field equations are a useful guide, but it is not expected that they accurately describe the behavior of the system in any finite dimension.
The average magnetization of a single spin at time $t$ is
$\langle\sigma^{\alpha}_{i}(t)\rangle=\sum_{\vec{\sigma}}\sigma^{\alpha}_{i}P({\vec{\sigma}},t).$ (18)
Using Eq. (13) and changing summation variables, we get
@@
Substituting the rates (Eq. (8)) and the selfish energy (Eq. (11)) we obtain
@@
where we have used the identity ${x\tanh{xb}=\tanh{b}}$ for $x=\pm 1$. Applying the mean-field approximation, $\langle g(\vec{\sigma})\rangle=g(\langle\vec{\sigma}\rangle)$ and defining the average magnetization as $\langle\sigma^{\alpha}_{i}\rangle\equiv m^{\alpha}_{i}$ yields
$\frac{{\rm d}m_{i}^{\alpha}}{{\rm d}t}=-\frac{m_{i}^{\alpha}}{\tau}+\frac{1}{\tau}\tanh\left(\frac{J\sum_{j\,{\rm nn\,of}\,i}m_{j}^{\alpha}+K_{\alpha\beta}m_{i}^{\beta}}{k_{B}T}\right).$ (21)
@@
where $D\equiv J/(k_{B}T)$ and $\tilde{K}_{\alpha\beta}\equiv K_{\alpha\beta}/(k_{B}T)$. Equation (23) is the mean-field equation of the general nonreciprocal Ising model. Note that it is invariant under the diagonal up-down $\mathbb{Z}_{2}$ symmetry
$\mathscr{S}_{1}:(m_{A},m_{B})\mapsto(-m_{A},-m_{B}).$ (24)
@@ V The fully anti-symmetric case
In this section as well as in Secs. VIIXI, we will study the fully anti-symmetric case, in which $K_{+}=0$. To simplify the notations, we write $K_{-}\equiv K$ (and $\tilde{K}_{-}\equiv\tilde{K}$). In this case, the mean-field equation reads
@@
where $\varepsilon_{\alpha\beta}$ is the Levi-Civita symbol. Equation (25) has the symmetry
$\mathscr{S}_{2}:(m_{A},m_{B})\mapsto(-m_{B},m_{A})$ (26)
in addition to that in Eq. (24), so overall it is invariant under the cyclic group $C_{4}$.
Fig. 3: Mean-field (MF) flow in $\left(m_{A},m_{B}\right)$ space. Arrows correspond to the vector $\vec{v}=(\partial_{t}m_{A},\partial_{t}m_{B})$ calculated from the spatially homogeneous full mean-field equation, Eq. (25). Fixed points are represented by dots. The limit cycle in the swap phase is highlighted by a semi-transparent blue line. Nonreciprocal inter-species coupling is $\tilde{K}=0.4$, and reciprocal coupling between same-species nearest neighbors is (a) $\tilde{J}=0.8$ , (b), $\tilde{J}=1.3$, and (c) $\tilde{J}=1.8$, showing the disorder, swap, and static-order phases, respectively.